40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jan 20 15:59:02 2012, kiwamu, Update, Green Locking, Y arm ALS noige budget Yarm_ALS_2012Jan19.png
    Reply  Fri Jan 20 16:24:50 2012, kiwamu, Update, Green Locking, Y arm ALS : time series time_series.png
       Reply  Fri Jan 20 17:05:59 2012, rana, Update, Green Locking, Y arm ALS : time series 
Message ID: 6214     Entry time: Fri Jan 20 15:59:02 2012     Reply to this: 6215
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Y arm ALS noige budget 

One of my goals in this week is : measurement of the current best ALS noise budget.

Last night I took a new noise spectra of the Y arm ALS, which is shown in the attached figure below.

The displacement of the arm cavity observed from the IR PDH is at 66 pm in rms. In the measurement the arm length was stabilized with the ALS technique.

 

Yarm_ALS_2012Jan19.png

 

ELOG V3.1.3-